OpenStudy (anonymous):

how do i make y = 2-(4/(X+1)) be in the form x= ... please help i got to y-2 = -(4/X+1)

OpenStudy (sosa4954):

wouldnt you want to distribute first?

OpenStudy (anonymous):

im not sure

OpenStudy (sosa4954):

yeah you would have to distribute -1 to everything

OpenStudy (anonymous):

okay how would i do that?

OpenStudy (anonymous):

im confused

OpenStudy (sosa4954):

take -1 and mutiple 4 with it you get....

OpenStudy (anonymous):

-4?

OpenStudy (sosa4954):

yes, and then you distribute again take -4 mutiply that with x and -4 with -1 whats the new equation?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

distribute what?

OpenStudy (sosa4954):

it would be y=2--4/-4x+4

OpenStudy (anonymous):

okay but how do i put it in terms of x = ...

OpenStudy (sosa4954):

+-4 to both sides

OpenStudy (sosa4954):

so it would be -4+y=2-4/4

OpenStudy (anonymous):

okay which would be y-4 = 2-1

OpenStudy (anonymous):

but where did x go?

OpenStudy (sosa4954):

Behined the 4 sorry